OpenStudy (anonymous):

((x-y)/2)+y, please explain how to get the answer

OpenStudy (anonymous):

What exactly are you looking for? The value of x?

OpenStudy (anonymous):

what is the question? what is the instruction on you book say? it's like saying this: what's the answer to x+y....

OpenStudy (anonymous):

Anyway, if you're looking for the value of x, it should be like: (x-y)/2 + y = 0 (x-y)/2 = -y Transpose y to the other side 2[(x-y)/2] = (-y)2 Multiply both sides by 2 and cancel the denominator 2 x-y = -2y x = -y Transpose -y to the other side, which makes it +y and add to -2y Now since x= -y, just substitute that one to the original equation (-y-y)/2 + y = 0 ^Not sure if that one's right. I didn't kinda understand your question. :))